OpenStudy (sloppycanada):

At the amusement park, you decide to ride the Ferris wheel, which has a maximum height of 50 meters and a diameter of 35 meters. It takes the wheel five minutes to make one revolution. Write the sinusoidal function, f(t), that models the height of your chair at any time, t. f(t) = 35sin(2pi/5t)+50?

OpenStudy (baru):

no...

OpenStudy (baru):

the question says, at the the highest point, the chair is 50m high. in your equation we get 35+50=85

OpenStudy (sloppycanada):

Shoot we do. And the max height is 50..

OpenStudy (baru):



OpenStudy (sloppycanada):

f(x) = 17.5sin(2pi/5t) + 32.5?

OpenStudy (baru):

yep :)

OpenStudy (sloppycanada):

At the amusement park, you decide to ride the Ferris wheel, which has a maximum height of 100 meters and a diameter of 50 meters. It takes the wheel seven minutes to make one revolution. Write the sinusoidal function, f(t), that models the height of your chair at any time, t. 25sin(2pi/7t) + 75?

OpenStudy (baru):

well inside the bracket its \(\frac{2\pi}{7}t\) when 't' is in minutes

OpenStudy (sloppycanada):

Right sorry. T should be outside the brackets. 25sin(2pi/7)t + 75

OpenStudy (baru):

\[25\sin( \frac{2 \pi}{7}t) +75\]